# Debiased Nonparametric Regression for Statistical Inference and Distributionally Robustness

### v2

@@ Abstract
This study proposes a debiasing method for smooth nonparametric estimators. While machine learning techniques such as random forests and neural networks have demonstrated strong predictive performance, their theoretical properties remain relatively underexplored. Specifically, many modern algorithms lack assurances of pointwise asymptotic normality and uniform convergence, which are critical for statistical inference and robustness under covariate shift and have been well-established for classical methods like Nadaraya-Watson regression. To address this, we introduce a model-free debiasing method that guarantees these properties for smooth estimators derived from any nonparametric regression approach. By adding a correction term that estimates the conditional expected residual of the original estimator, or equivalently, its estimation error, we obtain a debiased estimator with proven pointwise asymptotic normality, and uniform convergence. These properties enable statistical inference and enhance robustness to covariate shift, making the method broadly applicable to a wide range of nonparametric regression problems.
@@ 1 Introduction
This paper addresses the problem of nonparametric regression. Let $X\in{\mathcal{X}}$ be a covariate and $Y\in{\mathcal{Y}}$ a target variable, where ${\mathcal{X}}\in[0,1]^{d}$ and ${\mathcal{Y}}\subset{\mathbb{R}}$ represent the covariate and target spaces, respectively. While we primarily focus on the case where $X$ is one-dimensional for simplicity, the results presented in this paper extend to $d$-dimensional covariates.
Let $P$ denote the joint distribution of $(X,Y)$, and define the regression function under $P$ as
@@
where ${\mathbb{E}}_{P}\big{[}\cdot\big{]}$ represents the expectation operator with respect to $P$. We assume access to a dataset $\Big{\{}(X_{i},Y_{i})\Big{\}}_{i=1}^{n}$, where $(X_{i},Y_{i})$ are independent and identically distributed samples from a true distribution $P_{0}$. The true regression function is denoted by $f_{0}=f_{P_{0}}$. Our objective is to estimate $f_{0}$ based on the observations $\big{\{}(X_{i},Y_{i})\big{\}}_{i=1}^{n}$.
@@
We denote the distribution induced by a regression function $f$ as $P_{f}$, and the expectation operator under $P_{f}$ as ${\mathbb{E}}_{f}$. For a real-valued vector $z$, let $\|z\|$ denote its Euclidean norm. For a measurable function $h\colon{\mathcal{X}}\to{\mathbb{R}}$, let $\|h\|_{\infty}\coloneqq\sup_{x\in{\mathcal{X}}}|h(x)|$ denote the sup-norm, and $\|h\|_{2}\coloneqq\sqrt{{\mathbb{E}}\left[h(X)^{2}\right]}$ denote the $L^{2}$ norm. t
@@ 1.1 Content of this study
In Section 2, we define our debiased estimator in nonparametric regression. Our proposed estimator consists of the following three steps: estimate the regression function $f_{0}$ using any smooth nonparametric estimator; estimate the estimation error of the first stage estimator using the local polynomial regression; combine the first and second stage estimators. Here, we use the smoothness in the meaning that the either of the target nonparametric estimator, the true regression function, or their difference belongs to the Hölder class.
In Section 3, we show that our debiased nonparametric regression estimators have asymptotic normality and uniform convergence under mild conditions, like the smoothness. Furthermore, our estimator has the doubly robust property; that is, either of the first-stage nonparametric regression estimator or the second-stage conditional expected residual estimator is consistent, the resulting debiased estimator is consistent.
@@ 1.2 Related work
Doubly robust and debiased estimation has garnered significant attention across various fields, including statistics, economics, epidemiology, and machine learning. Doubly robust estimators are widely used in the context of semiparametric statistics and causal inference (van der Vaart, 1998; Bang & Robins, 2005). With the development of machine learning regression models, debiasing methods utilizing doubly robust estimators have also been extensively studied (Chernozhukov et al., 2018), particularly in conjunction with sample splitting techniques (Klaassen, 1987; van der Vaart, 2002; Zheng & van der Laan, 2011). While much of the existing literature focuses on parametric or semiparametric models, this study extends these approaches to nonparametric regression.
Nonparametric regression in machine learning faces several challenges:
  • For many machine learning estimators, the convergence rate to the true function is unknown, particularly with respect to minimax optimality.
  • Asymptotic normality remains unproven for many machine learning estimators, including random forests (Breiman, 2001).
@@
  • Uniform convergence is not established for many estimators, such as neural networks (Schmidt-Hieber & Zamolodtchikov, 2024).
  • Even for machine learning methods that theoretically satisfy these properties, verifying the required conditions can be challenging.
@@
Minimax optimality and asymptotic normality are crucial in statistical analysis, motivating efforts to establish these properties. For example, Wager & Athey (2018) and Mourtada et al. (2020) demonstrate asymptotic normality and minimax optimality for their modified random forests, respectively, but not for the original random forest proposed by Breiman (2001). These simplifications facilitate desirable theoretical results but may degrade empirical performance.
@@
Uniform convergence is particularly important for robustness under distributional shifts, such as covariate shift (Shimodaira, 2000), where nonparametric regression models are known to be effective. However, uniform convergence is rarely established for modern nonparametric regression estimators, including classical methods like local linear and series regression. This limitation stems from their data-adaptive nature and the prevalent use of empirical process arguments, which focus on population risk measures, such as mean squared error, rather than sup-norm bounds for uniform convergence. Schmidt-Hieber & Zamolodtchikov (2024) addresses this issue by demonstrating restricted uniform optimality for neural network regression in the one-dimensional covariate case, but their approach is not easily generalizable.
@@
Our method shares some motivations with debiasing approaches in high-dimensional regression problems. For instance, the Lasso estimator, while widely used, introduces a bias that diminishes with sample size at a slower rate than $\sqrt{n}$, hindering asymptotic normality (Tibshirani, 1996; Bühlmann & van de Geer, 2011). To mitigate this, approaches such as van de Geer et al. (2014) incorporate bias-correction terms, yielding improved asymptotic properties like normality and efficiency (Janková & van de Geer, 2018).
@@
This study generalizes our earlier work on doubly robust methods for nonparametric regression discontinuity design (Kato, 2024). Independently, Chernozhukov et al. (2024) develops conditional influence functions with similar motivations, building on work by Ichimura & Newey (2022).
From a technical perspective, this study is heavily influenced by Belloni et al. (2011) and Kennedy et al. (2024). Belloni et al. (2011) investigates various properties of nonparametric series estimators, which were subsequently utilized in Kennedy et al. (2024) to study minimax nonparametric estimators for conditional average treatment effects. We leverage these technical findings, alongside classical results from Stone (1980, 1982) and Tsybakov (2008).
@@ 2 The debiased estimator
We randomly split the observations ${\mathcal{D}}$ into two datasets, ${\mathcal{D}}^{(1)}$ and ${\mathcal{D}}^{(2)}$, satisfying ${\mathcal{D}}={\mathcal{D}}^{(1)}\cup{\mathcal{D}}^{(2)}$. For simplicity, let $n$ be an even value, and define $m=n/2$. Then, for each $\ell\in\{1,2\}$ we denote
@@
where ${\mathcal{I}}^{(\ell)}$ is the index set of ${\mathcal{D}}^{(\ell)}$; that is, ${\mathcal{I}}^{(1)}\cup{\mathcal{I}}^{(2)}=\{1,2,\dots,n\}$.
We consider the following three-stage estimation: for each point $x_{0}\in{\mathcal{X}}$ of interest:
@@
Estimate $f_{0}$ using any smooth model and denote the estimator by $\widehat{f}_{n}$.
@@
Estimate the conditional expected residual of the first-stage estimator ${\mathbb{E}}\big{[}Y-f(X)\mid X=x_{0}\big{]}$ or, equivalently, the estimation error $f_{0}(x_{0})-\widehat{f}_{n}(x_{0})$. We denote the estimator by
@@
Summing the first and second stage estimators and obtain
@@
as a debiased estimator of $f_{0}$.
As explained in Section 3, we can show the asymptotic normality and uniform convergence for the debiased estimator. In contrast, many of the modern machine learning methods lack those properties, even though their performances are guaranteed for the population risk.
In this section, we briefly explain the construction of the first-stage estimator $\widehat{f}_{n}$ and the second-stage estimator $\widehat{b}_{n}$.
@@ 2.1 First-stage nonparametric regression
For the first-stage estimator $\widehat{f}_{n}$, we can use any regression methods if $f_{0}-\widehat{f}_{n}$ is smooth. There can be several definitions for the smoothness. In this study, for simplicity, we focus on the smoothness in the meaning of the Hölder class (Definition 3.1). We note that both $f_{0}$ and $\widehat{f}_{n}$ do not have to be smooth if the difference $f_{0}-\widehat{f}_{n}$ is smooth.
We emphasize that in our analysis, we do not require specific properties for $\widehat{f}_{n}$ except for the smoothness. The first-stage estimator $\widehat{f}_{n}$ can converge to the true function $f_{0}$ with a very slow rate or even be inconsistent. Unless the smoothness conditions are satisfied, our asymptotic theoretical results hold, although the empirical finite-sample performance depends on the choice of the first stage estimator.
@@ 2.2 Second-stage conditional expected residual estimation
For each element $X_{j}$ of covariates $X=(X_{1}\ X_{2}\ \cdots\ X_{d})^{\top}\in{\mathcal{X}}$, let us define the Legendre polynomial series basis $\widetilde{\rho}\colon{\mathbb{R}}\to{\mathbb{R}}$ as
$\widetilde{\rho}(X_{j})\coloneqq\Big{(}\rho_{0}(X_{j})\ \ \rho_{1}(X_{j})\ \ \cdots\ \ \rho_{\lfloor\ell\rfloor}(X_{j})\Big{)}^{\top},$
where
$\widetilde{\rho}_{m}(X_{j})\coloneqq\sum^{m}_{k=0}(-1)^{\ell+m}\sqrt{2m+1}\begin{pmatrix}m\\ k\end{pmatrix}\begin{pmatrix}m+k\\ k\end{pmatrix}X^{k}_{j}.$
Then, we define $\rho(X_{i})$ to be the corresponding tensor product of all interactions of $\widetilde{\rho}_{m}(X_{1,i}),\dots,\widetilde{\rho}_{m}(X_{d,i})$.
By using the Legendre polynomial basis, we define the local linear regression as follows:
$\widehat{\beta}_{n}(x)\coloneqq\begin{pmatrix}\widehat{\beta}_{1,n}(x)&\widehat{\beta}_{2,n}(x)&\cdots&\widehat{\beta}_{\ell,n}(x)\end{pmatrix}^{\top}$
@@
where $K\colon{\mathcal{X}}\to{\mathbb{R}}$ is a kernel function, defined as
$K(x)\coloneqq K_{h}(x)\coloneqq\mathbbm{1}\left[\big{\|}x-x_{0}\big{\|}\leq h\right].$
Then, the conditional expected residual estimator corresponds to $\widehat{\beta}_{1,n}(x)$, the first element of $\widehat{\beta}_{n}(x)$; that is,
$\widehat{b}_{n}(x)=\widehat{\beta}_{1,n}(x).$
This second-stage estimator corresponds to the bias-correction term in the construction of semiparametric efficient estimators (Schuler & van der Laan, 2024). Such an estimator is referred to as the bias-correction or one-step estimator in the context of semiparametric statistics (van der Vaart, 2002).
2.3 Discussion
Our estimator (1) is closely related to the influence function in conditional mean estimation (Kennedy, 2023). This type of estimator is also connected to the Neyman orthogonal score (Chernozhukov et al., 2018), which is almost mathematically equivalent to the canonical gradient in the one-step bias corection (Schuler & van der Laan, 2024). Note that in conditional mean estimation, the construction of the influence function becomes complicated due to the conditioning. For the influence function including the parameter represented via the conditional expected value, see Ichimura & Newey (2022). Independently and simultaneously of us, Chernozhukov et al. (2024) investigates a related topic, with different interest and method.
@@ 3 Convergence analysis
This section presents convergence analysis of $\widetilde{f}_{n}$. First, we define the Hölder class as follows:
@@
Let ${\mathcal{T}}$ be an interval in ${\mathbb{R}}^{d}$ and let $s$ and $L$ be two positive numbers. Let $D^{\alpha}=\frac{\partial^{\alpha}}{\partial x_{1}^{\alpha_{1}}\dots\partial x_{d}^{\alpha_{d}}}$ be the partial derivative operator. The Hölder class $\Sigma(s,L)$ on ${\mathcal{T}}$ is defined as the set of $\ell=\lfloor s\rfloor$ times differentiable function $f\colon{\mathcal{T}}\to{\mathbb{R}}$ satisfying the following conditions:
  • Its derivatives up to order $\lfloor\ell\rfloor$ are bounded as
    $\left|D^{\alpha}f(x)\right|\leq C<\infty\qquad\forall x\in\mathcal{X},$
    for all $\alpha=(\alpha_{1},\dots,\alpha_{d})$ with $\sum_{j}\alpha_{j}\leq\ell$.
  • Its $\ell$-order derivatives satisfying the Lipschitz condition: there exists a constant $0<C\infty$ such that
    $\left|D^{\beta}f(x)-D^{\beta}f(x^{\prime})\right|\leq C\|x-x^{\prime}\|^{s-\ell}\qquad\forall x,x^{\prime}\in\mathcal{X}$
    for all $\beta=(\beta_{1},\dots,\beta_{d})$ with $\sum_{j}\beta_{j}=\ell$.
Define
^$\widehat{{\mathcal{B}}}_{n}(x_{0})$ $\coloneqq\frac{1}{mh^{d}_{n}}\sum_{i\in{\mathcal{I}}^{(2)}}\rho\left(\frac{X_{i}-x_{0}}{h_{n}}\right)\rho\left(\frac{X_{i}-x_{0}}{h_{n}}\right)^{\top}K\left(\frac{X_{i}-x_{0}}{h_{n}}\right),$
$\widehat{w}_{h}(x,x_{0})$ $\coloneqq\frac{1}{h^{d}_{n}}\rho\left(0\right)^{\top}\widehat{{\mathcal{B}}}_{n}(x_{0})^{-1}\rho\left(\frac{x-x_{0}}{h_{n}}\right)K\left(\frac{x-x_{0}}{h_{n}}\right).$
Then, the second-stage estimator is given as
$\widehat{b}_{n}(x_{0})=\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}Y_{i}-\widehat{f}_{n}(X_{i})\Big{)}\widehat{w}_{h}(X_{i},x_{0}).$

### v3

@@ Abstract
This study proposes a debiasing method for smooth nonparametric estimators. While machine learning techniques such as random forests and neural networks have demonstrated strong predictive performance, their theoretical properties remain relatively underexplored. In particular, many modern algorithms lack guarantees of pointwise and uniform risk convergence, as well as asymptotic normality. These properties are essential for statistical inference and robust estimation and have been well-established for classical methods such as Nadaraya-Watson regression. To ensure these properties for various nonparametric regression estimators, we introduce a model-free debiasing method. By incorporating a correction term that estimates the conditional expected residual of the original estimator, or equivalently, its estimation error, into the initial nonparametric regression estimator, we obtain a debiased estimator that satisfies pointwise and uniform risk convergence, along with asymptotic normality, under mild smoothness conditions. These properties facilitate statistical inference and enhance robustness to covariate shift, making the method broadly applicable to a wide range of nonparametric regression problems.
@@ 1 Introduction
This study investigates the problem of nonparametric regression. While modern machine learning methods enable the estimation of complex regression functions, they often lack the theoretical guarantees that traditional estimators possess, such as pointwise and uniform mean squared error (MSE) convergence, as well as asymptotic normality. Asymptotic normality is crucial for statistical inference, whereas uniform risk convergence plays an important role in prediction under covariate shift. In this study, we propose a debiasing method for regression estimators and show that, under mild smoothness conditions, they achieve asymptotic normality and uniform convergence.
Here, we formulate the problem. Let $X\in{\mathcal{X}}$ be a $d$-dimensional covariate and $Y\in{\mathcal{Y}}$ be a target variable, where ${\mathcal{X}}\subseteq{\mathbb{R}}^{d}$ and ${\mathcal{Y}}\subset{\mathbb{R}}$ represent the covariate and target spaces, respectively. Let $P$ denote the joint distribution of $(X,Y)$, and define the regression function under $P$ as
@@
where ${\mathbb{E}}_{P}\big{[}\cdot\big{]}$ represents the expectation operator with respect to $P$. We assume access to observations $\big{\{}(X_{i},Y_{i})\big{\}}_{i=1}^{n}$, where $(X_{i},Y_{i})$ are independent and identically distributed samples from a true distribution $P_{0}$. The true regression function is denoted by $f_{0}=f_{P_{0}}$. Our objective is to estimate $f_{0}$ from the observations $\big{\{}(X_{i},Y_{i})\big{\}}_{i=1}^{n}$, aiming for pointwise and uniform risk convergence, as well as asymptotic normality.
@@
For a real-valued vector $z$, let $\|z\|$ denote its Euclidean norm. For a measurable function $h\colon{\mathcal{X}}\to{\mathbb{R}}$, let $\|h\|_{\infty}\coloneqq\sup_{x\in{\mathcal{X}}}|h(x)|$ denote the sup-norm, and $\|h\|_{2}\coloneqq\sqrt{{\mathbb{E}}\left[h(X)^{2}\right]}$ denote the $L^{2}$ norm.
@@ 1.1 Content of this study
In Section 2, we define our debiased estimator for nonparametric regression. Our proposed estimator consists of the following three steps: (1) estimate the regression function $f_{0}$ using a nonparametric regression method; (2) estimate the expected conditional residual, or equivalently, the estimation error of the first-stage estimator, using local polynomial regression; and (3) add the second-stage estimator to the first-stage estimator.
In Section 3, we present an example of our proposed debiased estimator, applying local polynomial regression for conditional expected residual estimation.
In Section 4, we establish that our debiased nonparametric regression estimator achieves pointwise and uniform MSE convergence if either (a) the first-stage nonparametric regression estimator and the true regression function, or (b) their difference, belongs to the Hölder class. Using pointwise MSE convergence, we also prove asymptotic normality. Furthermore, our estimator possesses the doubly robust property: if either the first-stage nonparametric regression estimator or the second-stage conditional expected residual estimator is consistent, then the resulting debiased estimator is also consistent.
The key takeaway is that our debiased estimator ensures desirable theoretical properties, typically associated with classical methods, for a wide range of modern first-stage estimators. As long as certain smoothness conditions are met, our method effectively corrects bias in a broad class of initial estimators, equipping the resulting estimator with strong theoretical guarantees.
@@ 1.2 Related work
Doubly robust and debiased estimation has attracted significant attention across various fields, including statistics (Tsiatis, 2007), economics (Chernozhukov et al., 2018), epidemiology (Bang & Robins, 2005), and machine learning (Kallus & Uehara, 2020). With the advancement of machine learning regression models, debiasing methods based on doubly robust estimators have been extensively studied (Chernozhukov et al., 2018), particularly in conjunction with sample splitting techniques (Klaassen, 1987; van der Vaart, 2002; Zheng & van der Laan, 2011). While much of the existing literature focuses on parametric or semiparametric models, this study extends these approaches to nonparametric regression. The application of debiased machine learning to nonparametric regression is discussed in Kennedy (2023), though their focus differs from ours.
Nonparametric regression in machine learning presents several challenges:
@@
  • Pointwise and uniform MSE convergence is not established for many estimators, such as neural networks and random forests.
  • Asymptotic normality remains unproven for numerous machine learning estimators.
@@
Pointwise and uniform MSE convergence and asymptotic normality are critical in statistical analysis, motivating significant efforts to establish these properties. For instance, Wager & Athey (2018) and Mourtada et al. (2020) demonstrate asymptotic normality and optimality, respectively, for their modified random forests, but not for the original random forest of Breiman (2001).
@@
Pointwise and uniform convergence is particularly relevant for robustness under distributional shifts, such as covariate shift (Shimodaira, 2000). These convergence properties are rarely established for modern nonparametric estimators, whereas classical methods like local linear and series regression possess them. This gap arises in part due to the data-adaptive nature of modern methods and the frequent use of empirical process arguments, which often focus on population MSE, where expectations are taken over the covariate distribution rather than on pointwise or sup-norm MSE. Schmidt-Hieber & Zamolodtchikov (2024) addresses this concern by demonstrating restricted uniform optimality for neural network regression in the one-dimensional covariate setting, but this approach does not generalize easily.
@@
Our method shares some motivations with debiasing approaches in high-dimensional regression. For example, the Lasso estimator, while widely used, introduces a bias that decreases with sample size at a slower rate than $\sqrt{n}$, thereby hindering asymptotic normality (Tibshirani, 1996; Bühlmann & van de Geer, 2011). To mitigate this, van de Geer et al. (2014) proposes incorporating bias-correction terms, leading to improved asymptotic properties such as normality and efficiency (Janková & van de Geer, 2018).
@@
This study generalizes our previous work on doubly robust methods for nonparametric regression discontinuity design (Kato, 2024). Concurrently with our work, Chernozhukov et al. (2024) and van der Laan et al. (2025) also explore double machine learning in the context of nonparametric regression, though their emphases differ substantially from ours. For example, Chernozhukov et al. (2024) develop conditional influence functions, which are related to the framework of Ichimura & Newey (2022).
@@ 2 The debiased estimator
In this section, we define our debiased estimator in a general way. First, we randomly split the observations ${\mathcal{D}}$ into two datasets, ${\mathcal{D}}^{(1)}$ and ${\mathcal{D}}^{(2)}$, such that ${\mathcal{D}}={\mathcal{D}}^{(1)}\cup{\mathcal{D}}^{(2)}$. For simplicity, let $n$ be even and set $m=n/2$. Then, for each $\ell\in\{1,2\}$, we define
@@
where ${\mathcal{I}}^{(\ell)}$ is the index set of ${\mathcal{D}}^{(\ell)}$, i.e., ${\mathcal{I}}^{(1)}\cup{\mathcal{I}}^{(2)}=\{1,2,\dots,n\}$.
We consider the following three-stage estimation for each point $x_{0}\in{\mathcal{X}}$ of interest:
@@
Estimate $f_{0}$ using any smooth model, and denote this estimator by $\widehat{f}_{n}$.
@@
Estimate the conditional expected residual of the first-stage estimator, ${\mathbb{E}}\big{[}Y-f(X)\mid X=x_{0}\big{]}$, or equivalently, the estimation error $f_{0}(x_{0})-\widehat{f}_{n}(x_{0})$. Denote this estimator by $\widehat{b}_{n}(x_{0})$.
@@
Sum the first- and second-stage estimators to obtain
@@
which serves as a debiased estimator of $f_{0}$.
The term debiased estimator refers to our objective of correcting bias arising from the first-stage estimator. In many nonparametric and machine learning methods, bias can be problematic. By subtracting the estimated residual, ${\mathbb{E}}\big{[}Y-\widehat{f}_{n}(X)\mid X=x_{0}\big{]}$, we compensate for the difference between $f_{0}$ and $\widehat{f}_{n}$, resulting in an estimator with improved convergence and inferential properties.
Our estimator (1) is closely related to the influence function in conditional mean estimation (Kennedy, 2023). It is also connected to the Neyman orthogonal score (Chernozhukov et al., 2018), which is essentially equivalent to the canonical gradient in one-step bias correction (Schuler & van der Laan, 2024). Notably, in conditional mean estimation, constructing the influence function is more intricate due to the conditioning. For an influence function that involves parameters represented via the conditional expected value, see Ichimura & Newey (2022). Independently and concurrently, Chernozhukov et al. (2024) investigates a related topic with different motivations and methods.
3 The debiased nonparametric regression with local polynomial conditional expected residual estimation
In this section, we present an example of our debiased estimator. We employ local polynomial regression to estimate the conditional expected residual. For simplicity, we focus on the case where the covariate is one-dimensional ($d=1$) and ${\mathcal{X}}=[0,1]$.
@@ 3.1 First-stage nonparametric regression
For the first-stage estimator $\widehat{f}_{n}$, any regression method can be used, provided that smoothness conditions are met. As we show later, our key assumption is that the difference $f_{0}-\widehat{f}_{n}$ is smooth. While various notions of smoothness exist, for simplicity, we focus on smoothness in the sense of the Hölder class (Definition 4.1). Notably, neither $f_{0}$ nor $\widehat{f}_{n}$ needs to be smooth individually, as long as their difference $f_{0}-\widehat{f}_{n}$ is smooth.
We emphasize that our analysis does not impose any specific requirements on $\widehat{f}_{n}$ beyond smoothness. The first-stage estimator $\widehat{f}_{n}$ may converge to $f_{0}$ at an arbitrarily slow rate or even be inconsistent. Nevertheless, as long as the smoothness conditions hold, our asymptotic theoretical results remain valid. However, to achieve strong finite-sample performance, a high-quality first-stage estimator is typically preferred. In practice, methods such as random forests, neural networks, or other advanced machine learning techniques can be employed, provided they satisfy the smoothness conditions to a reasonable extent.
@@ 3.2 Second-stage conditional expected residual estimation
In the second-stage estimation of the conditional expected residual, we employ local polynomial regression. We define the $\ell$-th order polynomial basis as follows:
$\rho(u)\coloneqq\begin{pmatrix}1&u&u^{2}/2!&\cdots&u^{\ell}/\ell!\end{pmatrix}^{\top}.$
In the theoretical analysis, the value of $\ell$ is chosen based on the smoothness of $f_{0}-\widehat{f}_{n}$.
Using this polynomial basis, we define the local polynomial regression estimator as follows:
$\widehat{\beta}_{n}(x)\coloneqq\begin{pmatrix}\widehat{\beta}_{0,n}(x)&\widehat{\beta}_{1,n}(x)&\widehat{\beta}_{2,n}(x)&\cdots&\widehat{\beta}_{\ell,n}(x)\end{pmatrix}^{\top}$
@@
where $K\colon{\mathcal{X}}\to{\mathbb{R}}$ is a kernel function defined by
$K(u)\coloneqq K_{h}(u)\coloneqq\mathbbm{1}\left[\big{|}u\big{|}\leq h\right].$
The conditional expected residual estimator is then given by the first component of $\widehat{\beta}_{n}(x)$:
$\widehat{b}_{n}(x)=\widehat{\beta}_{0,n}(x).$
@@ 4 Convergence analysis
This section presents a convergence analysis of $\widetilde{f}_{n}$ defined in Section 3. We begin by defining the Hölder class.
@@
Given an interval ${\mathcal{T}}\subset{\mathbb{R}}$, and positive constants $\beta$ and $L$, the Hölder class $\Sigma(\beta,L)$ on ${\mathcal{T}}$ is the set of $\ell=\lfloor\beta\rfloor$ times differentiable functions $f\colon{\mathcal{T}}\to{\mathbb{R}}$ whose $\ell$-times derivative $f^{(\ell)}$ satisfies
$\Big{|}f^{(\ell)}(x)-f^{(\ell)}(z)\Big{|}\leq L\big{|}x-z\big{|}^{\beta-\ell}\quad\forall x,z\in{\mathcal{T}}.$
4.1 Closed-form solution
The debiased estimator with local polynomial regression admits a closed-form solution. We define
^$\widehat{{\mathcal{B}}}_{n}(x_{0})$ $\coloneqq\frac{1}{mh_{n}}\sum_{i\in{\mathcal{I}}^{(2)}}\rho\left(\frac{X_{i}-x_{0}}{h_{n}}\right)\rho\left(\frac{X_{i}-x_{0}}{h_{n}}\right)^{\top}K\left(\frac{X_{i}-x_{0}}{h_{n}}\right),$
$\widehat{w}_{h}(x,x_{0})$ $\coloneqq\frac{1}{h_{n}}\rho(0)^{\top}\widehat{{\mathcal{B}}}_{n}(x_{0})^{-1}\rho\left(\frac{x-x_{0}}{h_{n}}\right)K\left(\frac{x-x_{0}}{h_{n}}\right).$
Then, the second-stage estimator is given by
$\widehat{b}_{n}(x_{0})=\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\left(Y_{i}-\widehat{f}_{n}(X_{i})\right)\widehat{w}_{h}(X_{i},x_{0}).$
Thus, the debiased estimator is expressed as
$\widetilde{f}_{n}(x_{0})=\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\left(Y_{i}-\widehat{f}_{n}(X_{i})\right)\widehat{w}_{h}(X_{i},x_{0})+\widehat{f}_{n}(x_{0}).$